OpenStudy (wwe123):

if a,b,c are in GP,prove that 1/a,1/b,1/c are also in GP

OpenStudy (raden):

if a,b,c are in GP, so b/a = c/b or b/c = a/b now, if 1/a,1/b,1/c are also in GP, so must be satisfies 1/b * a = 1/c * b or a/b = b/c. proof

OpenStudy (wwe123):

Gp = geometric progression

OpenStudy (anonymous):

Generally speaking if a^n = b^(n+k) then the inverse of each side reveals 1/a^n = 1/b^(n+k) Or (1/a)^n = (1/b)^(n+k) If they are right next to each other in progression then k = 1.

OpenStudy (wwe123):

i am not understand @RadEn

OpenStudy (wwe123):

i am not understand

OpenStudy (raden):

each ratio of geo series is always same right ?

OpenStudy (wwe123):

yes and

OpenStudy (raden):

if given a,b,c, so the ratio (r) = b/a = c/b right ?

OpenStudy (wwe123):

right and b=ar & c=ar^2 but then

OpenStudy (raden):

i dont think b=ar and c=ar^2 but, from b/a = c/b, it means also b/c = a/b ?

OpenStudy (wwe123):

\[\frac{ b }{ a }=\frac{ c }{ b }=r\]

OpenStudy (raden):

b/a = c/b ---> b * b = a * c ----> b/c = a/b, agree ?

OpenStudy (wwe123):

yes

OpenStudy (raden):

now, do same idea for series of 1/a, 1/b, 1/c r = (1/b)/(1/a) = (1/c)/(1/b) simplify and u must show to get b/c = a/b too

OpenStudy (usukidoll):

what! ummm...I'm not even there when it comes to proofs. I'm just a beginner.

OpenStudy (wwe123):

thanks to all.................